If (1+x+2x²)^10=a′+a1x+a2x²+.....+a20x^20,then a′+a2+a4+........a20 is equal to 

Grade:

1 Answers

Arun
5 years ago
When x = 1
 
Then ao + a1 +a2+ a3+ a4+........+a20 = 4^10
 
When x = -1
 
ao -a1+a2-a3 +a4- ............+a20 = 2^10
 
Now add both the equations
 
(ao + a2 + a4+ a6+......+ a20) = 1/2 (4^10 + 2^10)
 
 
Hope it helps